# Optimal Switching in Networked Control Systems: Finite Horizon

### 29 Mar 2026 v1

@@
Optimal Switching in Networked Control Systems: Finite Horizon
@@ 1 Abstract
In this work, we first prove that the separation principle holds for switched LQR problems under i.i.d. zero-mean disturbances with a symmetric distribution. We then solve the dynamic programming problem and show that the optimal switching policy is a symmetric threshold rule on the accumulated disturbance since the most recent update, while the optimal controller is a discounted linear feedback law independent of the switching policy.
@@ 2 Introduction
As the name suggests, NCS research investigates control loops in which one or more feedback paths traverse a communication network. Early contributions introduced packet losses and delays into feedback control and examined the resulting performance and stability. For example, Åström and Bernhardsson compared the mean variance of a continuous‑time system under periodic sampling and state‑dependent (event‑triggered) sampling, assuming impulsive control actions at sampling instants [6]. Tabuada proposed an event‑triggered feedback policy for linear time‑invariant (LTI) systems, where the maximum update period depends on the estimation error; the controller employs a zero‑order hold (ZOH), updating only at sampling times and holding the control value constant in between [36]. Montestruque and Antsaklis investigated the stochastic stability of NCSs with time‑varying update intervals (independent and identically distributed (i.i.d.) or Markov‑chain driven) modeling the sampled system as a jump‑linear system [28]. Heemels et al. introduced self‑triggered sampling, where each update time is computed at the preceding update and compared with event‑triggered policies [14]. Zhang et al. separately examined the effects of delays and packet dropouts on stability. Delays were handled by employing system estimates in both full‑state and partial‑state feedback, while packet dropouts were analyzed under periodic transmissions to determine the minimum stable transmission rate [43]. Finally, Lyapunov-based analyses by Heemels and colleagues established the maximum allowable transmission interval (MATI) and maximum allowable delay (MAD) that ensure NCS stability [15].
@@
Fundamentally, any NCS consists of two key design components. These are the switching [^1] In the literature, this mechanism is referred to by various terms such as scheduling, sampling, transmission, or switching. Throughout this paper, we adopt the term switching. mechanism and the controller (see Fig. 1). In this work, we focus on their joint optimization. We model system uncertainties as i.i.d. zero-mean disturbances with a symmetric distribution and employ a standard quadratic cost on state deviations and control effort. This formulation is known as the Linear Quadratic Regulator (LQR), since the objective is to regulate the system state to the origin. This criterion is widely used in the NCS literature [29].
@@
Our aim is to find the optimal switching and control policies for the finite-horizon LQR problem under an average switching-rate constraint. The main difficulty arises from the dual effect, whereby transmission decisions influence future estimation errors while control actions affect the evolution of the estimator. This coupling between the switching mechanism and the controller breaks the classical separation principle, which ordinarily allows the estimator and controller to be designed independently without loss of optimality [8]. To regain separability, much of the existing literature imposes simplifying assumptions, most commonly by restricting attention to a subclass of policies under which the conditional mean of the accumulated disturbance is zero. We define such policies as symmetric policies. While this assumption renders the problem analytically tractable, it restricts the policy space, raising the question of whether this subclass is optimal. Crucially, however, the optimality of this restricted policy space has yet to be proven, representing a significant gap in the current understanding of switched LQR problems. In what follows, we review how this joint optimization problem has been treated in the literature, highlighting the widespread use of the symmetry assumption, sometimes imposed explicitly by assuming zero-mean disturbances and sometimes adopted implicitly through Bernoulli sampling models, or symmetric threshold rules.
@@
In [24], the switching-rate constraint was relaxed through a Lagrangian formulation, and separation was enforced by assuming zero-mean disturbances. The design was then recast as an equivalent optimal-estimation problem to enable numerical solutions. An equivalent form of the symmetry assumption was adopted implicitly in [10], where the design was reformulated as an age-of-information (AoI) minimization problem and a heuristic scheduler was proposed. In a two-hop multi-system setting [7], disturbances were likewise constrained to be zero-mean, and AoI- and value-of-information (VoI)-based schedulers were compared numerically. Building on this line of work, a VoI metric was introduced in [40], for which a threshold policy was proposed, and the concept was further refined in [35] with corresponding sampling thresholds.
@@
Other studies have focused on the joint switch and controller design under symmetric threshold policies. In [31], a threshold was imposed on the squared estimation error, under which the linear feedback law was found to be optimal. The relation between this threshold and the Lagrange multiplier was investigated in [27], with convergence and stability properties analyzed. Energy constraints at the sensor were incorporated in [19], where a threshold based on the conditional covariance of the estimation error was proposed. More recent work [2] employed a Q-learning–based reinforcement-learning algorithm under the same symmetric assumption, while varying delays were addressed in [22] through a mixed-integer control problem solved offline with Riccati methods and integer programming. From a differential-game perspective [1], continuous and intermittent sensing players were contrasted, leading to AoI-based threshold rules. Finally, the disturbance-driven estimation dynamics were correctly modeled in [21], where a relaxed dynamic program for the sampler was formulated and an approximate solution proposed due to computational complexity.
@@
  • We revisit the constrained joint design of switching and control policies for scalar LTI systems, aiming to clarify subtle underlying assumptions.
  • Most importantly, we prove that symmetric policies are optimal for the finite-horizon switched LQR problem under a switching-rate constraint, thereby establishing that the separation principle holds.
  • Finally, we solve the associated dynamic programming (DP) problem and derive the optimal switching and control policies, which take the form of a symmetric threshold rule on the estimation error and a discounted linear feedback law that is independent of the switching policy, respectively.
We begin by analyzing the information structure of the system and identifying when the classical separation principle breaks down. We then discuss why at least a unit-delay must be introduced between switch and controller. Then, we introduce the class of symmetric policies and demonstrate that under these policies, the dual effect is neutralized, allowing the estimation and control tasks to decouple. We rigorously prove that this subclass is not only tractable but optimal for the finite-horizon LQR problem under a switching-rate constraint. Finally, we derive the closed-form solution for the optimal system, which consists of a linear feedback control law and a symmetric threshold switching policy based on the accumulated disturbances.
@@ 3 System Model
where $X_{k}$ denotes the system state, $U_{k}$ the control input, and $W_{k}$ an additive disturbance. The disturbance sequence $\{W_{k}\}$ is assumed to be i.i.d., zero-mean, with finite variance $\sigma_{W}^{2}$, and symmetric around zero. The controller receives intermittent information through a switch with a fixed $\tau\geq 1$ delay, so that the observation available to the controller is
@@
where $D_{k}\in\{0,1\}$ denotes the switching decision at time $k$. In the sequel, we discuss why at least a one-step delay must be introduced between the switch and the controller.
Let $t_{k}$ denote the most recent switching time prior to $k$, defined as
@@
The information vector at the switch plays a key role in the optimal‐switching problem. Various approaches and information sets appear in the literature. In the most general form, we denote the switch’s information by
@@
Remark 1.
Under delay-free feedback, the controller information set includes $Y_{k}=D_{k}X_{k}\in I_{k}^{\mathcal{C}}$. Note that $Y_{k}$ depends explicitly on the switching decision $D_{k}$. However, at time $k$, the switch only has access to $D_{k-1}\in I_{k}^{\mathcal{S}}$. Since the controller is located downstream of the switch, including $Y_{k}=D_{k}X_{k}$ in the controller information set violates causality [41]. Therefore, at least a one-step delay must be introduced between the switch and the controller to ensure a causal information structure.
Note that several works in the literature assume delay-free feedback and study separation or optimal switching in the switched LQR problem [24, 27, 25, 31, 7, 30, 19, 18]. However, as shown in Remark 1, such formulations are non-causal, rendering the associated DP recursion ill-posed.
The control and switching policies are defined as
@@
where the sets $\Gamma$ and $\Pi$ contain all admissible control and switching decision rules that satisfy the usual measurability and integrability conditions. We write $\boldsymbol{\mu}=(\gamma_{0},\gamma_{1},\dots)\in\Gamma$ and $\boldsymbol{\pi}=(f_{0},f_{1},\dots)\in\Pi$ for generic control- and switching-policy sequences, respectively. Because the two sequences are designed jointly, $\boldsymbol{\mu}$ and $\boldsymbol{\pi}$ can be coupled.
@@
Consider the system (1)–(4) and the optimization problem $P_{1}$. The optimal controller is given by the linear feedback law  [^2] We assume a sufficiently large horizon length so that the steady-state Riccati solution is applicable. The extension to the transient case is straightforward.
@@
Given the optimal controller in (9), the objective function in (7) can be reformulated in terms of the estimation error $\mathcal{E}_{k}$. The goal is then to determine the optimal switching policy that minimizes this reformulated cost.
@@
In formulation $P_{2}$, the cost depends only on the estimation error. Consequently, if the estimation‐error dynamics can be written independently of the control inputs, the control and switching problems decouple, i.e., the separation principle holds. Note that without the separation principle, although Theorem 1 gives the optimal control law $U_{k}=-L\,\hat{X}_{k},$ the state estimate $\hat{X}_{k}$ itself depends on the switching policy $\boldsymbol{\pi}$. Until $\boldsymbol{\pi}$ is fixed, the estimator remains undefined and the closed‐loop law cannot be completed. Only once $\boldsymbol{\pi}$ is chosen can both the estimator and the corresponding optimal controller be fully specified.
@@ 4 Symmetric Policies
We demonstrate that the statistical coupling between switching decisions and disturbances fundamentally prevents estimation error decoupling unless this symmetry is imposed. Ultimately, this section provides the necessary foundation for Section 5, where we move beyond tractability to rigorously prove that these symmetric policies are, in fact, optimal.
@@ 4.1 Estimation Error Dynamics and the Role of Symmetry
For $\tau\geq 1$ and the switching instant $t_{k}$ , we obtain
@@
Step (a) holds because $X_{t_{k}}$ is included in $I^{\mathcal{C}}_{t_{k}+\tau}$, so its conditional error vanishes. Step (b) follows because, for $j=0,\dots,\tau-1$, the noise variables $W_{t_{k}+j}$ are zero-mean and independent of the sigma-algebra $\mathcal{F}_{k,\tau}:=\sigma\!\bigl(I^{\mathcal{C}}_{t_{k}+\tau}\bigr)$, which contains only information available up to time $t_{k}$ (including the switching decision $D_{t_{k}}$) but not the realizations of $W_{t_{k}},\ldots,W_{t_{k}+\tau-1}$. Hence, $\mathbb{E}[\,W_{t_{k}+j}\mid\mathcal{F}_{k,\tau}]=0,$.
@@
there exist admissible switching policies under which
@@
For all $\tau\geq 1$, we define the subclass of admissible switching policies as
@@
This subclass enforces that conditioning on future switching decisions does not alter the conditional mean of the state. Additionally, note that for policies that do not belong to $\Pi^{\prime}$, the following strict inequality holds:
@@
there exist admissible switching policies under which the disturbance $W_{k}$ becomes statistically dependent on the switching decision. Consequently, the conditional mean of the disturbance may be nonzero, i.e.,
@@
We define the class of switching policies for which the accumulated disturbance remains conditionally independent of the controller’s information over multiple steps as
@@
Intuitively, these are referred to as symmetric policies because, under such policies, the conditional distribution of the accumulated disturbance remains symmetric (zero mean) with respect to the controller’s information, irrespective of the switching decisions. In contrast, non-symmetric policies may induce a bias in the conditional expectation of the accumulated disturbance.
@@
Observe that $P_{3}$ depends only on the estimation error dynamics and is independent of the control input. Consequently, the control and switching policies decouple, allowing us to focus exclusively on the switching policy and thereby invoke the separation principle. However, since not all switching policies satisfy the error evolution in (26), it remains to show that symmetric policies are optimal. Establishing this result renders problems $P_{2}$ and $P_{3}$ equivalent, in other words, separation holds.

### 02 May 2026 v3

# Separation is Optimal for LQR under Intermittent Feedback

@@
Separation is Optimal for LQR under Intermittent Feedback
@@ 1 Abstract
In this work, we first prove that the separation principle holds for communication-constrained LQR problems under i.i.d. zero-mean disturbances with a symmetric distribution. We then solve the dynamic programming problem and show that the optimal scheduling policy is a symmetric threshold rule on the accumulated disturbance since the most recent update, while the optimal controller is a discounted linear feedback law independent of the scheduling policy.
@@ 2 Introduction
As the name suggests, NCS research investigates control loops in which one or more feedback paths traverse a communication network. Early contributions introduced packet losses and delays into feedback control and examined the resulting performance and stability. For example, Åström and Bernhardsson compared the mean variance of a continuous‑time system under periodic sampling and state‑dependent (event‑triggered) sampling, assuming impulsive control actions at sampling instants [6]. Tabuada proposed an event‑triggered feedback policy for linear time‑invariant (LTI) systems, where the maximum inter-update period depends on the estimation error; the controller employs a zero‑order hold (ZOH), updating only at sampling times and holding the control value constant in between [36]. Montestruque and Antsaklis investigated the stochastic stability of NCSs with time‑varying update intervals (independent and identically distributed (i.i.d.) or Markov‑chain driven) modeling the sampled system as a jump‑linear system [28]. Heemels et al. introduced self‑triggered sampling, where each update time is computed at the preceding update and compared with event‑triggered policies [14]. Zhang et al. separately examined the effects of delays and packet dropouts on stability. Delays were handled by employing system estimates in both full‑state and partial‑state feedback, while packet dropouts were analyzed under periodic transmissions to determine the minimum stable transmission rate [43]. Finally, Lyapunov-based analyses by Heemels and colleagues established the maximum allowable transmission interval (MATI) and maximum allowable delay (MAD) that ensure NCS stability [15].
@@
Fundamentally, any NCS consists of two key design components. These are the scheduling [^1] In the literature, this mechanism is referred to by various terms such as scheduling, sampling, transmission, or switching. Throughout this paper, we adopt the term scheduling. mechanism and the controller (see Fig. 1). In this work, we focus on their joint optimization. We model system uncertainties as i.i.d. zero-mean disturbances with a symmetric distribution and employ a standard quadratic cost on state deviations and control effort. This formulation is known as the Linear Quadratic Regulator (LQR), since the objective is to regulate the system state to the origin. This criterion is widely used in the NCS literature [29].
@@
Our aim is to find the optimal scheduling and control policies for the finite-horizon LQR problem under an average communication-rate constraint. The main difficulty arises from the dual effect, whereby transmission decisions influence future estimation errors while control actions affect the evolution of the estimator. This coupling between the scheduling mechanism and the controller breaks the classical separation principle, which ordinarily allows the estimator and controller to be designed independently without loss of optimality [8]. To regain separability, much of the existing literature imposes simplifying assumptions, most commonly by restricting attention to a subclass of policies under which the conditional mean of the accumulated disturbance is zero. We define such policies as symmetric policies. While this assumption renders the problem analytically tractable, it restricts the policy space, raising the question of whether this subclass is optimal. Crucially, however, the optimality of this restricted policy space has yet to be proven, representing a significant gap in the current understanding of communication-constrained LQR problems. Next, we review how this joint optimization problem has been treated in the literature, highlighting the widespread use of the symmetry assumption, sometimes imposed explicitly by assuming zero-mean disturbances and sometimes adopted implicitly through Bernoulli sampling models, or symmetric threshold rules.
@@
In [24], the communication-rate constraint was relaxed through a Lagrangian formulation, and separation was enforced by assuming zero-mean disturbances. The design was then recast as an equivalent optimal-estimation problem to enable numerical solutions. An equivalent form of the symmetry assumption was adopted implicitly in [10], where the design was reformulated as an age-of-information (AoI) minimization problem and a heuristic scheduler was proposed. In a two-hop multi-system setting [7], disturbances were likewise constrained to be zero-mean, and AoI- and value-of-information (VoI)-based schedulers were compared numerically. Building on this line of work, a VoI metric was introduced in [40], for which a threshold policy was proposed, and the concept was further refined in [35] with corresponding sampling thresholds.
@@
Other studies have focused on the joint scheduler and controller design under symmetric threshold policies. In [31], a threshold was imposed on the squared estimation error, under which the linear feedback law was found to be optimal. The relation between this threshold and the Lagrange multiplier was investigated in [27], with convergence and stability properties analyzed. Energy constraints at the sensor were incorporated in [19], where a threshold based on the conditional covariance of the estimation error was proposed. More recent work [2] employed a Q-learning–based reinforcement-learning algorithm under the same symmetric assumption, while varying delays were addressed in [22] through a mixed-integer control problem solved offline with Riccati methods and integer programming. From a differential-game perspective [1], continuous and intermittent sensing players were contrasted, leading to AoI-based threshold rules. Finally, the disturbance-driven estimation dynamics were correctly modeled in [21], where a relaxed dynamic program for the sampler was formulated and an approximate solution proposed due to computational complexity.
@@
  • We revisit the constrained joint design of scheduling and control policies for scalar LTI systems, aiming to clarify subtle underlying assumptions.
  • Most importantly, we prove that symmetric policies are optimal for the finite-horizon LQR problem under a communication-rate constraint, thereby establishing that the separation principle holds.
  • Finally, we solve the associated dynamic programming (DP) problem and derive the optimal scheduling and control policies, which take the form of a symmetric threshold rule on the estimation error and a discounted linear feedback law that is independent of the scheduling policy, respectively.
We begin by analyzing the information structure of the system and identifying when the classical separation principle breaks down. We then discuss why at least a unit-delay must be introduced between the scheduling mechanism and controller. Then, we introduce the class of symmetric policies and demonstrate that under these policies, the dual effect is neutralized, allowing the estimation and control tasks to decouple. We rigorously prove that this subclass is not only tractable but optimal for the finite-horizon LQR problem under a communication-rate constraint. Finally, we derive the closed-form solution for the optimal system, which consists of a linear feedback control law and a symmetric threshold scheduling policy based on the accumulated disturbances.
@@ 3 System Model
where $X_{k}$ denotes the system state, $U_{k}$ the control input, and $W_{k}$ an additive disturbance. The disturbance sequence $\{W_{k}\}$ is assumed to be i.i.d., zero-mean, with finite variance $\sigma_{W}^{2}$, and symmetrically distributed about zero (not necessarily Gaussian). The controller receives intermittent information through the scheduling mechanism with a fixed $\tau\geq 1$ delay, so that the observation available to the controller is
@@
where $D_{k}\in\{0,1\}$ denotes the scheduling decision at time $k$. In the sequel, we discuss why at least a one-step delay must be introduced between the scheduling mechanism and controller.
Let $t_{k}$ denote the most recent update time prior to $k$, defined as
@@
The information vector at the scheduler plays a key role in the optimal scheduling problem. Various approaches and information sets appear in the literature. In the most general form, we denote the scheduler information by
@@
Causality of the Information Structure:
A policy is said to be causal if, at time $k$, the control input $U_{k}$ depends only on the information available up to time $k$, i.e., $U_{k}$ must be measurable with respect to the controller information set $I_{k}^{\mathcal{C}}$ defined in (3). In delay-free formulations, the controller information set $I_{k}^{\mathcal{C}}$ is often taken to include the measurement $Y_{k}=D_{k}X_{k}$. However, from (4), the scheduling decision $D_{k}$ is generated by the scheduler at time $k$ based on $I_{k}^{\mathcal{S}}$, which includes $X_{k}$ but not $D_{k}$ itself prior to its realization. Since the controller is located downstream of the scheduler, the value of $D_{k}$ becomes available to the controller only after the scheduler makes its decision.
Therefore, if $U_{k}$ is required to be generated at time $k$ based on $I_{k}^{\mathcal{C}}$, including $Y_{k}=D_{k}X_{k}$ in $I_{k}^{\mathcal{C}}$ implicitly assumes that the controller has instantaneous access to $D_{k}$. This creates a circular dependence, as $U_{k}$ would depend on a variable that is generated at the same time instant and not yet available at the moment of decision, violating causality in the sense of [41].
To ensure a well-defined causal information structure, the controller must rely only on variables available strictly prior to or at the decision time. This can be enforced by introducing at least a one-step delay between the scheduling mechanism and controller, so that the controller uses $D_{k-1}$ instead of $D_{k}$ when computing $U_{k}$.
Note that several works in the literature assume delay-free feedback and study separation or optimal scheduling in the communication-constrained LQR problem [24, 27, 25, 31, 7, 30, 19, 18]. As discussed, such formulations implicitly rely on information patterns that violates causality, which can affect the validity of the associated dynamic programming formulations.
With the causal information structure now specified, we define the control and scheduling policies as
@@
where the sets $\Gamma$ and $\Pi$ contain all admissible control and scheduling decision rules that satisfy the usual measurability and integrability conditions. We write $\boldsymbol{\mu}=(\gamma_{0},\gamma_{1},\dots)\in\Gamma$ and $\boldsymbol{\pi}=(f_{0},f_{1},\dots)\in\Pi$ for generic control- and scheduling-policy sequences, respectively. Because the two sequences are designed jointly, $\boldsymbol{\mu}$ and $\boldsymbol{\pi}$ can be coupled.
@@
Consider the system (1)–(4) and the optimization problem $P_{1}$. The optimal controller is given by the linear feedback law
@@
Given the optimal controller in (9), the objective function in (7) can be reformulated in terms of the estimation error $\mathcal{E}_{k}$. The goal is then to determine the optimal scheduling policy that minimizes this reformulated cost.
@@
Note that in Theorem 1, we use the steady-state solution of the Riccati equation for notational simplicity, which leads to a simplified form of the equivalent problem $P_{2}$. In the transient case, the only difference is the appearance of the term $L_{k}(r+b^{2}P_{k+1})$ in the summation. This term does not fundamentally affect the analysis. For completeness, we provide the key equations that change under the transient controller in Appendix H.
In formulation $P_{2}$, the cost depends only on the estimation error. Consequently, if the estimation‐error dynamics can be written independently of the control inputs, the control and scheduling problems decouple, i.e., the separation principle holds. Note that without the separation principle, although Theorem 1 gives the optimal control law $U_{k}=-L\,\hat{X}_{k},$ the state estimate $\hat{X}_{k}$ itself depends on the scheduling policy $\boldsymbol{\pi}$. Until $\boldsymbol{\pi}$ is fixed, the estimator remains undefined and the closed‐loop law cannot be completed. Only once $\boldsymbol{\pi}$ is chosen can both the estimator and the corresponding optimal controller be fully specified.
@@ 4 Symmetric Policies
We demonstrate that the statistical coupling between scheduling decisions and disturbances fundamentally prevents estimation error decoupling unless this symmetry is imposed. Ultimately, this section provides the necessary foundation for Section 5, where we move beyond tractability to rigorously prove that these symmetric policies are, in fact, optimal.
@@ 4.1 Estimation Error Dynamics and the Role of Symmetry
For $\tau\geq 1$ and the scheduling instant $t_{k}$ , we obtain
@@
Step (a) holds because $X_{t_{k}}$ is included in $I^{\mathcal{C}}_{t_{k}+\tau}$, so its conditional error vanishes. Step (b) follows because, for $j=0,\dots,\tau-1$, the noise variables $W_{t_{k}+j}$ are zero-mean and independent of the sigma-algebra $\mathcal{F}_{k,\tau}:=\sigma\!\bigl(I^{\mathcal{C}}_{t_{k}+\tau}\bigr)$, which contains only information available up to time $t_{k}$ (including the scheduling decision $D_{t_{k}}$) but not the realizations of $W_{t_{k}},\ldots,W_{t_{k}+\tau-1}$. Hence, $\mathbb{E}[\,W_{t_{k}+j}\mid\mathcal{F}_{k,\tau}]=0,$.
@@
there exist admissible scheduling policies under which
@@
In much of the existing literature, this dependence is implicitly neglected by assuming that (15) holds with equality, which simplifies the analysis. However, this assumption is not valid in general and requires justification.
@@
For all $\tau\geq 1$, we define the subclass of admissible scheduling policies as
@@
This subclass enforces that conditioning on future scheduling decisions does not alter the conditional mean of the state. Additionally, note that for policies that do not belong to $\Pi^{\prime}$, the following strict inequality holds:
@@
there exist admissible scheduling policies under which the disturbance $W_{k}$ becomes statistically dependent on the scheduling decision. For instance, if the scheduler employs a threshold-based policy of the form $D_{k}=\mathds{1}_{\{W_{k-1}>\beta\}}$ for some $\beta>0$, then the event $D_{k}$ directly reveals information about the disturbance $W_{k-1}$, inducing statistical dependence. Consequently, the conditional mean of the disturbance may be nonzero, i.e.,
@@
We define the class of scheduling policies for which the accumulated disturbance remains conditionally independent of the controller’s information over multiple steps as
@@
Intuitively, these are referred to as symmetric policies because, under such policies, the conditional distribution of the accumulated disturbance remains symmetric (zero mean) with respect to the controller’s information, irrespective of the scheduling decisions. In contrast, non-symmetric policies may induce a bias in the conditional expectation of the accumulated disturbance.
@@
Observe that $P_{3}$ depends only on the estimation error dynamics and is independent of the control input. Consequently, the control and scheduling policies decouple, allowing us to focus exclusively on the scheduling policy and thereby invoke the separation principle. However, since not all scheduling policies satisfy the error evolution in (26), it remains to show that symmetric policies are optimal. Establishing this result renders problems $P_{2}$ and $P_{3}$ equivalent, in other words, separation holds.